# How Effective are Large Time Series Models in Hydrology? A Study on Water Level Forecasting in Everglades

### v1

@@ Abstract
The Everglades play a crucial role in flood and drought regulation, water resource planning, and ecosystem management in the surrounding regions. However, traditional physics-based and statistical methods for predicting water levels often face significant challenges, including high computational costs and limited adaptability to diverse or unforeseen conditions. Recent advancements in large time series models have demonstrated the potential to address these limitations, with state-of-the-art deep learning and foundation models achieving remarkable success in time series forecasting across various domains. Despite this progress, their application to critical environmental systems, such as the Everglades, remains underexplored. In this study, we fill the gap by investigating twelve task-specific models and five time series foundation models across six categories for a real-world application focused on water level prediction in the Everglades. Our primary results show that the foundation model Chronos significantly outperforms all other models while the remaining foundation models exhibit relatively poor performance. Moreover, the performance of task-specific models varies with the model architectures. Lastly, we discuss the possible reasons for the varying performance of models.
@@ 1 Introduction
The Everglades is a unique ecosystem characterized by vast, swampy grassland, especially containing sawgrass and seasonally covered by slowly moving water. Water levels in the Everglades play a pivotal role in the prevention of floods and droughts, sustainable water resource planning, recreational and infrastructure management and ecosystem management Paudel et al. (2020); Gibble et al. (2020); Saberski et al. (2022). For example, the U.S. Everglades National Park in South Florida, spanning 2,357 square miles, as shown in Figure 1, has a profound influence on urban development, flood control, agricultural expansion, and natural hydrology. Therefore, accurate water level forecasting in the Everglades is of significance, promoting environmental sustainability and benefiting society.
@@
Water levels are typically recorded as time series, and recent advancements in deep learning (DL) have demonstrated its significant potential in time series forecasting. These models excel at capturing nonlinear dependencies across time and variables, enabling both short- and long-term predictions. For example, recurrent neural networks (RNNs) and convolutional neural networks (CNNs) have been used to learn the temporal dependencies for the prediction of water level Shi et al. (2023b). Graph neural networks, often integrated with sequential models, have been utilized to learn complex spatiotemporal relationships from multiple environmental variables Kazadi et al. (2022). More broadly, state-of-the-art models for time series forecasting have demonstrated remarkable performance across diverse domains Liu and Wang (2024), such as electricity, weather, traffic, healthcare, and energy. Most of them leverage Transformer-based architectures Vaswani (2017), which excel at capturing long-term correlations among elements within time series. However, these models usually require high computation costs due to the self-attention scheme with $\mathcal{O}(L^{2})$ time and memory complexity, where $L$ is the sequence length. Although several variants of Transformers Li et al. (2019); Zhou et al. (2021); Nie et al. (2022) have improved the computational complexity to $\mathcal{O}(L\log L)$ or $\mathcal{O}(L)$, significant challenges still persist. Furthermore, some researchers have questioned whether the self-attention mechanism captures temporal information, claiming that even linear models with $\mathcal{O}(L)$ complexity can outperform Transformer-based models for long-term time series forecasting Zeng et al. (2023). Additionally, some researchers have attempted to leverage large language models (LLMs) for time series forecasting Jin et al. (2023). These efforts focus on adapting the comprehensive prior knowledge of LLMs to the time series domain Jin et al. (2024). While these models have demonstrated exceptional performance, they are fully task-specific, requiring retraining from scratch for each new dataset or experimental setting.
@@
Inspired by the success of large language models (LLMs) in natural language processing, researchers have explored the development of foundation models tailored to process time series data. These models adopt a training paradigm similar to that used for LLMs, where each time point in a time series is treated as analogous to a “word” or “token” in a sentence. By leveraging the self-supervised learning technique, these models are pre-trained on large-scale time series datasets to capture complex temporal patterns. The resulting pre-trained models can be fine-tuned to adapt to various downstream tasks. The first time series foundation model is called TimeGPT Garza and Mergenthaler (2023). Other examples include TimesFM Das et al. (2023), Timer Liu et al. (2024a), Moirai Woo et al. (2024) and Chronos Ansari et al. (2024), etc. They have demonstrated remarkable performance in zero-shot setting, reducing or eliminating the need for task-specific retraining.
@@
While these task-specific and foundation models have shown unparalleled performance in time series forecasting across various domains, their application to hydrology is largely unexplored. Here we study the performance of these models for water level prediction in the Everglades, with profound implications to hydrology. Our contributions include:
@@ 2.1 Region Selection and Data Collection
The study domain focuses on the Everglades National Park (see Figure 2), including various hydrological structures and stations, and measurements, such as water levels, water flows, precipitation, and evapotranspiration values. We downloaded the corresponding data from the Everglades Depth Estimation Network (EDEN) Haider et al. (2020) and South Florida Water Management District’s Environmental Database (DBHYDRO) South Florida Water Management District (2024). The selection process resulted in a dataset with 37 variables spanning daily records of water levels, water flows, rainfall, and potential evapotranspiration (PET). We provide a summary of the data in Table 3 (see Appendix A).
Figure 2: Study area encompassing portion of the Everglades National Park, covering key hydrological structures and stations.
@@ 2.2 State-of-The-Art Models
2.2.1 Linear-based Models
@@
DLinear and NLinear, introduced in Zeng et al. (2023), are two simple yet effective linear models designed for long-term time series forecasting. DLinear decomposes the raw input data into two components: a trend component, extracted using a moving average kernel, and a remainder (seasonal) component. Each component is processed independently by a single-layer linear model to learn its feature representation, and the two components are then summed up to generate the final prediction. NLinear operates by first subtracting the last value of the input sequence from all elements in the sequence. The adjusted input is then passed through a linear layer to learn the feature representation, after which the subtracted value is added back to generate the final prediction.
2.2.2 MLP-based Models
@@
NBEATS Oreshkin et al. (2019) employs a deep neural architecture based on backward and forward residual links and a deep stack of fully-connected layers. The stacked architecture comprises multiple fully connected blocks, enabling iterative refinement of predictions through backward and forward forecasting. TimeMixer Wang et al. (2024) is based on multi-layer perceptrons (MLPs) and designed for multi-scale time series modeling, capturing both microscopic seasonal and macroscopic trends. TSMixer Chen et al. (2023) is a novel architecture comprising stacked MLPs, which learns from both time and feature dimensions with mixing operations. TSMixerx Chen et al. (2023) is an extension of TSMixer, allowing additional exogenous features, including static features and future information, such as product categories and promotional events.
2.2.3 Transformer-based Models
@@
PatchTST Nie et al. (2022) segments long time series into multiple non-overlapping patches of fixed length, aggregating time steps into subseries-level patches to capture rich local information that is otherwise unavailable at the point level. Additionally, it leverages the Transformer architecture with self-attention mechanisms to model temporal dependencies across these patches, enabling the capture of long-term dependencies while reducing computational overhead. Informer Zhou et al. (2021) addresses the computational challenges of vanilla Transformers in long-horizon time series forecasting by introducing three key innovations. First, it employs a self-attention mechanism, which reduces the time and memory complexity from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L\log L)$ by focusing on the most critical query-key pairs. Second, it incorporates a self-attention distillation process, enabling the model to retain essential temporal dependencies while effectively handling long input sequences. iTransformer Liu et al. (2023) modifies the traditional Transformer architecture by implementing the attention mechanism across multiple variables rather than individual time points, where each time series is embedded as a token. This model has shown promising performance in capturing cross-variable dependencies in multivariate time series data.
2.2.4 KAN-based Models
@@
KANs (Kolmogorov-Arnold Networks) Liu et al. (2024b) are an alternative to MLPs, using the Kolmogorov-Arnold approximation theorem, where splines are learned to approximate more complex functions. Unlike the MLP, the nonlinear functions are learned at the edges. At the nodes, the edge functions are merely summed up. RMoK (Reversible Mixture of KAN) Han et al. (2024b) builds on the KAN concept and employs a mixture-of-experts approach to capture both global and local patterns in time series data. It uses a gating network to assign variables to specialized KAN experts, such as Wav-KAN, JacobiKAN, and TaylorKAN. These experts utilize different mathematical functions and polynomials to effectively model complex time series relationships.
2.2.5 LLM-based Models
@@
TimeLLM Jin et al. (2023) is a reprogramming framework to repurpose LLMs without changes for general time series forecasting. First, it reprograms the input time series with text prototypes by Prompt-as-Prefix (PaP), and then both time series and prompts are fed into a native LLM to align the two modalities. In other words, TimeLLM transforms a forecasting task into a “language task” that can be tackled by an off-the-shelf LLM.
2.2.6 Foundation Models
@@
TimeGPT Garza and Mergenthaler (2023) is the first foundation model pre-trained with enormous amounts of time series data. It is capable of generating accurate zero-shot inference for diverse datasets not seen during training. It is engineered to process time series of different characteristics while accommodating different input sizes and forecasting horizons. TimesFM Das et al. (2023) is a time series foundation model that adopts a decoder-style attention mechanism for pretraining. It processes time series data by segmenting it into patches, enabling TimesFM to capture both local and global temporal dependencies. Timer Liu et al. (2024a) curates large-scale datasets with up to 1 billion time points, unifying heterogeneous time series into a single series into single-series sequence (S3) format, thus developing a GPT-style architecture for large time series models (LTSMs). Moirai Woo et al. (2024) extends the TimesFM architecture to the any-variate setting by “flattening” multivariate time series to a long single sequence. It achieves universal time series forecasting by training a masked encoder-based Transformer to address three challenges inherent in time series data: i) multiple frequencies, ii) multivariate forecasting, and iii) varying distributions. Chronos Ansari et al. (2024) tokenizes time series values using scaling and quantization into a fixed vocabulary and trains the transformer-based architecture via a cross-entropy loss function.
@@ 2.3 Forecasting Task
where $\mathcal{F}_{\theta}$ represents a predictive model with learnable parameters $\theta$. In what follows, we will drop the superscripts and use $\textbf{X}_{t-w+1:t}$ and $\textbf{Y}_{t+1:t+k}$.
@@
Models Lead Time NP205 P33 G620 NESRS1 NESRS2 Overall
@@
NLinear Zeng et al. (2023) Linear-based 7 0.195 0.528 0.114 0.508 0.086 0.132 0.081 0.176 0.063 0.113 0.108 0.291
@@
DLinear Zeng et al. (2023) Linear-based 7 0.165 0.245 0.081 0.117 0.094 0.132 0.071 0.118 0.062 0.110 0.095 0.145
@@
NBEATS Oreshkin et al. (2019) MLP-based 7 0.149 0.223 0.062 0.106 0.064 0.116 0.056 0.107 0.049 0.099 0.076 0.130
@@
TimeMixer Wang et al. (2024) MLP-based 7 0.170 0.290 0.079 0.121 0.070 0.126 0.069 0.119 0.061 0.113 0.090 0.154
@@
TSMixer Chen et al. (2023) MLP-based 7 0.145 0.226 0.072 0.111 0.080 0.126 0.063 0.108 0.055 0.100 0.083 0.134
@@
TSMixerx Chen et al. (2023) MLP-based 7 0.366 0.570 0.154 0.205 0.171 0.234 0.139 0.211 0.129 0.204 0.192 0.285
@@
Informer Zhou et al. (2021) Transformer-based 7 0.279 0.389 0.134 0.174 0.160 0.207 0.155 0.193 0.155 0.195 0.177 0.232
@@
PatchTST Nie et al. (2022) Transformer-based 7 0.132 0.208 0.067 0.103 0.071 0.113 0.060 0.105 0.052 0.097 0.076 0.125
@@
iTransformer Liu et al. (2023) Transformer-based 7 0.205 0.320 0.080 0.120 0.083 0.135 0.072 0.118 0.070 0.111 0.102 0.161
@@
RMoK Han et al. (2024b) KAN-based Model 7 0.171 0.247 0.088 0.118 0.100 0.137 0.080 0.116 0.070 0.106 0.102 0.145
@@
TimeLLM Jin et al. (2023) LLM-based Model 7 0.228 0.308 0.106 0.135 0.125 0.160 0.090 0.124 0.080 0.121 0.126 0.170
@@
TimeGPT Garza and Mergenthaler (2023) Foundational Model 7 0.160 0.262 0.082 0.138 0.086 0.153 0.074 0.136 0.065 0.131 0.093 0.164
@@
TimesFM Das et al. (2023) Foundational Model 7 0.224 0.337 0.118 0.166 0.121 0.186 0.108 0.171 0.094 0.156 0.133 0.203
@@
Timer Liu et al. (2024a) Foundational Model 7 0.293 0.374 0.160 0.194 0.181 0.224 0.146 0.184 0.131 0.171 0.182 0.229
@@
Moirai Woo et al. (2024) Foundational Model 7 0.471 0.742 0.205 0.299 0.285 0.424 0.186 0.184 0.161 0.231 0.261 0.389
@@
Chronos Ansari et al. (2024) Foundational Model 7 0.094 0.132 0.047 0.059 0.042 0.054 0.036 0.051 0.028 0.045 0.049 0.075
@@
Table 1: Model performance for 5 stations (NP205, P33, G620, NESRS1, and NESRS2) for lead times of 7, 14, 21, and 28 days. The first 12 models are task-specific while the last 5 are pre-trained foundation models for time series modeling. The best results are in bold.

### v2

@@ Abstract
The Everglades play a crucial role in flood and drought regulation, water resource planning, and ecosystem management in the surrounding regions. However, traditional physics-based and statistical methods for predicting water levels often face significant challenges, including high computational costs and limited adaptability to diverse or unforeseen conditions. Recent advancements in large time series models have demonstrated the potential to address these limitations, with state-of-the-art deep learning and foundation models achieving remarkable success in time series forecasting across various domains. Despite this progress, their application to critical environmental systems, such as the Everglades, remains underexplored. In this study, we fill the gap by investigating twelve task-specific models and five time series foundation models across six categories for a real-world application focused on water level prediction in the Everglades. Our primary results show that the foundation model Chronos significantly outperforms all other models while the remaining foundation models exhibit relatively poor performance. We also noticed that the performance of task-specific models varies with the model architectures, and discussed the possible reasons. We hope our study and findings will inspire the community to explore the applicability of large time series models in hydrological applications. The code and data are available at https://github.com/rahuul2992000/Everglades-Benchmark.
@@ 1 Introduction
The Everglades is a unique ecosystem characterized by vast, swampy grassland, especially containing sawgrass, and seasonally covered by slowly moving water. Water levels in the Everglades play a pivotal role in the prevention of floods and droughts, sustainable water resource planning, recreational and infrastructure management, and ecosystem management (Paudel et al. 2020; Saberski et al. 2022). For example, the U.S. Everglades National Park in South Florida, spanning 2,357 square miles, as shown in Figure 1, has a profound influence on urban development, flood control, agricultural expansion, and natural hydrology. Therefore, accurate water level forecasting in the Everglades is vital, promoting environmental sustainability and benefiting society.
@@
Water levels are typically recorded as time series, and recent advancements in deep learning (DL) have demonstrated its significant potential in time series forecasting. These models excel at capturing nonlinear dependencies across time and variables, enabling both short- and long-term predictions (Kumar et al. 2023). For example, recurrent neural networks (RNNs) and convolutional neural networks (CNNs) have been used to learn the temporal dependencies for the prediction of water level (Shi et al. 2023b). Graph neural networks, often integrated with sequential models, have been utilized to learn complex spatiotemporal relationships from multiple environmental variables (Kazadi et al. 2022).. More broadly, state-of-the-art models for time series forecasting have demonstrated remarkable performance across diverse domains (Liu and Wang 2024), such as electricity, weather, traffic, healthcare, and energy. Most of them leverage Transformer-based architectures (Vaswani 2017), which excel at capturing long-term correlations among elements within time series. However, these models usually require high computation costs due to the self-attention scheme with $\mathcal{O}(L^{2})$ time and memory complexity, where $L$ is the sequence length. Although several variants of Transformers (Li et al. 2019; Zhou et al. 2021; Nie et al. 2022) have improved the computational complexity to $\mathcal{O}(L\log L)$ or $\mathcal{O}(L)$, significant challenges still persist. Furthermore, some researchers have questioned whether the self-attention mechanism captures temporal information, claiming that even linear models with $\mathcal{O}(L)$ complexity can outperform Transformer-based models for long-term time series forecasting (Zeng et al. 2023). Additionally, some researchers have attempted to leverage large language models (LLMs) for time series forecasting (Jin et al. 2023). These efforts focus on adapting the comprehensive prior knowledge of LLMs to the time series domain (Jin et al. 2024). While these models have demonstrated exceptional performance, they are fully task-specific, requiring retraining from scratch for each new dataset or experimental setting.
@@
Inspired by the success of large language models (LLMs) in natural language processing, researchers have explored the development of foundation models tailored to process time series data. These models adopt a training paradigm similar to that used for LLMs, where each time point in a time series is treated as analogous to a “word” or “token” in a sentence. By leveraging the self-supervised learning technique, these models are pre-trained on large-scale time series datasets to capture complex temporal patterns. The resulting pre-trained models can be fine-tuned to adapt to various downstream tasks. The first time series foundation model is called TimeGPT (Garza and Mergenthaler 2023). Other examples include TimesFM (Das et al. 2023), Timer (Liu et al. 2024a), Moirai (Woo et al. 2024) and Chronos (Ansari et al. 2024), etc. They have demonstrated remarkable performance in a zero-shot setting, reducing or eliminating the need for task-specific retraining.
@@
While these task-specific and foundation models have shown unparalleled performance in time series forecasting across various domains, their application to hydrology is largely unexplored. We study the applicability of these models for water level prediction in the Everglades, with profound implications in hydrology. Our contributions are:
@@ 2.1 Region Selection and Data Collection
The study domain focuses on the Everglades National Park (see Figure 2), including various hydrological structures and stations, and measurements, such as water levels, water flows, precipitation, and evapotranspiration values. We downloaded the corresponding data from the Everglades Depth Estimation Network (EDEN) (Haider et al. 2020) and South Florida Water Management District’s Environmental Database (DBHYDRO) (South Florida Water Management District 2024). The selection process resulted in a dataset with 37 variables spanning daily records of water levels, water flows, rainfall, and potential evapotranspiration (PET). We provide a summary of the data in Appendix A.
Figure 2: Study area: a portion of the Everglades National Park, covering key hydrological structures and stations.
@@ 2.2 State-of-The-Art Models
Linear-based Models
@@
DLinear and NLinear, introduced in (Zeng et al. 2023), are two simple yet effective linear models designed for long-term time series forecasting. DLinear decomposes the raw input data into two components: a trend component, extracted using a moving average kernel, and a remainder (seasonal) component. Each component is processed independently by a single-layer linear model to learn its feature representation, and the two components are then summed up to generate the final prediction. NLinear operates by first subtracting the last value of the input sequence from all elements in the sequence. The adjusted input is then passed through a linear layer to learn the feature representation, after which the subtracted value is added back to generate the final prediction.
MLP-based Models
@@
NBEATS (Oreshkin et al. 2019) employs a deep neural architecture based on backward and forward residual links and a deep stack of fully-connected layers. The stacked architecture comprises multiple fully connected blocks, enabling iterative refinement of predictions through backward and forward forecasting. TimeMixer (Wang et al. 2024) is based on multi-layer perceptrons (MLPs) and designed for multi-scale time series modeling, capturing both microscopic seasonal and macroscopic trends. TSMixer is a novel architecture comprising stacked MLPs, which learns from both time and feature dimensions with mixing operations. TSMixerx (Chen et al. 2023) is an extension of TSMixer, allowing additional exogenous features, including static features and future information, such as product categories and promotional events.
Transformer-based Models
@@
PatchTST (Nie et al. 2022) segments long time series into multiple non-overlapping patches of fixed length, aggregating time steps into subseries-level patches to capture rich local information that is otherwise unavailable at the point level. Additionally, it leverages the Transformer architecture with self-attention mechanisms to model temporal dependencies across these patches, enabling the capture of long-term dependencies while reducing computational overhead. Informer (Zhou et al. 2021) addresses the computational challenges of vanilla Transformers in long-horizon time series forecasting by introducing three key innovations. First, it employs a self-attention mechanism, which reduces the time and memory complexity from $\mathcal{O}(L^{2})$ to $\mathcal{O}(L\log L)$ by focusing on the most critical query-key pairs. Second, it incorporates a self-attention distillation process, enabling the model to retain essential temporal dependencies while effectively handling long input sequences. iTransformer (Liu et al. 2023) modifies the traditional Transformer architecture by implementing the attention mechanism across multiple variables rather than individual time points, where each time series is embedded as a token. This model has shown promising performance in capturing cross-variable dependencies in multivariate time series data.
KAN-based Models
@@
KANs (Kolmogorov-Arnold Networks) (Liu et al. 2024b) are an alternative to MLPs, using the Kolmogorov-Arnold approximation theorem, where splines are learned to approximate more complex functions. Unlike the MLP, the nonlinear functions are learned at the edges. At the nodes, the edge functions are merely summed up. RMoK (Reversible Mixture of KAN) (Han et al. 2024c) builds on the KAN concept and employs a mixture-of-experts approach to capture both global and local patterns in time series data. It uses a gating network to assign variables to specialized KAN experts, such as Wav-KAN, JacobiKAN, and TaylorKAN. These experts utilize different mathematical functions and polynomials to effectively model complex time series relationships.
LLM-based Models
@@
TimeLLM (Jin et al. 2023) is a reprogramming framework to repurpose LLMs without changes for general time series forecasting. First, it reprograms the input time series with text prototypes by Prompt-as-Prefix (PaP), and then both time series and prompts are fed into a native LLM to align the two modalities. In other words, TimeLLM transforms a forecasting task into a “language task” that can be tackled by an off-the-shelf LLM.
Foundation Models
@@
TimeGPT (Garza and Mergenthaler 2023) is the first foundation model pre-trained with enormous amounts of time series data. It is capable of generating accurate zero-shot inference for diverse datasets not seen during training. It is engineered to process time series of different characteristics while accommodating different input sizes and forecasting horizons. TimesFM (Das et al. 2023) is a time series foundation model that adopts a decoder-style attention mechanism for pretraining. It processes time series data by segmenting it into patches, enabling TimesFM to capture both local and global temporal dependencies. Timer (Liu et al. 2024a) curates large-scale datasets with up to 1 billion time points, unifying heterogeneous time series into a single series into single-series sequence (S3) format, thus developing a GPT-style architecture for large time series models (LTSMs). Moirai (Woo et al. 2024) extends the TimesFM architecture to the any-variate setting by “flattening” multivariate time series to a long single sequence. It achieves universal time series forecasting by training a masked encoder-based Transformer to address three challenges inherent in time series data: i) multiple frequencies, ii) multivariate forecasting, and iii) varying distributions. Chronos (Ansari et al. 2024) tokenizes time series values using scaling and quantization into a fixed vocabulary and trains the transformer-based architecture via a cross-entropy loss function.
@@ 2.3 Forecasting Task
where $\mathcal{F}_{\theta}$ represents a predictive model with learnable parameters $\theta$. The superscripts allall and waterwater represent all (e.g., water levels, flows, and other measurements) and water variables, respectively. In what follows, we will drop the superscripts and use $\textbf{X}_{t-w+1:t}$ and $\textbf{Y}_{t+1:t+k}$.
@@
Models
Lead
Time
NP205 P33 G620 NESRS1 NESRS2 Overall
@@
NLinear
(Zeng et al. 2023)
Linear-based
7 0.195 0.528 0.114 0.508 0.086 0.132 0.081 0.176 0.063 0.113 0.108 0.291
@@
DLinear
(Zeng et al. 2023)
Linear-based
7 0.165 0.245 0.081 0.117 0.094 0.132 0.071 0.118 0.062 0.110 0.095 0.145
@@
NBEATS
(Oreshkin et al. 2019)
MLP-based
7 0.149 0.223 0.062 0.106 0.064 0.116 0.056 0.107 0.049 0.099 0.076 0.130
@@
TimeMixer
(Wang et al. 2024)
MLP-based
7 0.170 0.290 0.079 0.121 0.070 0.126 0.069 0.119 0.061 0.113 0.090 0.154
@@
TSMixer
(Chen et al. 2023)
MLP-based
7 0.145 0.226 0.072 0.111 0.080 0.126 0.063 0.108 0.055 0.100 0.083 0.134
@@
TSMixerx
(Chen et al. 2023)
MLP-based
7 0.366 0.570 0.154 0.205 0.171 0.234 0.139 0.211 0.129 0.204 0.192 0.285
@@
Informer
(Zhou et al. 2021)
Transformer-based
7 0.279 0.389 0.134 0.174 0.160 0.207 0.155 0.193 0.155 0.195 0.177 0.232
@@
PatchTST
(Nie et al. 2022)
Transformer-based
7 0.132 0.208 0.067 0.103 0.071 0.113 0.060 0.105 0.052 0.097 0.076 0.125
@@
iTransformer
(Liu et al. 2023)
Transformer-based
7 0.205 0.320 0.080 0.120 0.083 0.135 0.072 0.118 0.070 0.111 0.102 0.161
@@
RMoK
(Han et al. 2024c)
KAN-based Model
7 0.171 0.247 0.088 0.118 0.100 0.137 0.080 0.116 0.070 0.106 0.102 0.145
@@
TimeLLM
(Jin et al. 2023)
LLM-based Model
7 0.228 0.308 0.106 0.135 0.125 0.160 0.090 0.124 0.080 0.121 0.126 0.170
@@
TimeGPT
(Garza and Mergenthaler 2023)
Foundational Model
7 0.160 0.262 0.082 0.138 0.086 0.153 0.074 0.136 0.065 0.131 0.093 0.164
@@
TimesFM
(Das et al. 2023)
Foundational Model
7 0.224 0.337 0.118 0.166 0.121 0.186 0.108 0.171 0.094 0.156 0.133 0.203
@@
Timer
(Liu et al. 2024a)
Foundational Model
7 0.293 0.374 0.160 0.194 0.181 0.224 0.146 0.184 0.131 0.171 0.182 0.229
@@
Moirai
(Woo et al. 2024)
Foundational Model
7 0.471 0.742 0.205 0.299 0.285 0.424 0.186 0.184 0.161 0.231 0.261 0.389
@@
Chronos
(Ansari et al. 2024)
Foundational Model
7 0.094 0.132 0.047 0.059 0.042 0.054 0.036 0.051 0.028 0.045 0.049 0.075
@@
Table 1: Performance across 5 stations (NP205, P33, G620, NESRS1, NESRS2) for lead times of 7, 14, 21, and 28 days. The first 12 models are task-specific, while the last 5 are pre-trained foundation models for time series. The best results are in bold.